OpenStudy (anonymous):

HELLLLLPPP SOS The following lines are parallel. 2x + 2y = 4 and y = -x + 5

OpenStudy (anonymous):

TRUE OR FALSE?

OpenStudy (anonymous):

you know the slope of each? that is what you need

OpenStudy (anonymous):

like y=mx+b? @satellite73

OpenStudy (anonymous):

yeah the second one looks like \(y=mx+b\) for sure, what is \(m\)?

OpenStudy (anonymous):

-1?@satellite73

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

for the slope of \[2x+2y=4\] you can solve for \(y\) in two steps to get the slope

OpenStudy (anonymous):

if you do not know how i can show you

OpenStudy (anonymous):

yeah can you show me please @satellite73

OpenStudy (anonymous):

subtract \(2x\)then divide by \(2\) \[2x+2y=4\\ 2y=-2x+4\\ y=-x+2\]

OpenStudy (anonymous):

so they are parallel because they have the same slope since m in this problem also is -1? @satellite73

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

thank you!! @satellite73